OpenStudy (anonymous):

simplify -x^3+2x(x-x^2)

OpenStudy (anonymous):

pretty much like the last one distribute the \(2x\) first then combine like terms

OpenStudy (anonymous):

need help doing it?

OpenStudy (anonymous):

would i distribute it to the numbers in the parenthesis?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

let me know what you get

OpenStudy (anonymous):

so the 2x(x-x^2) would be 2x^2-x^2? so that'd just be 2? bc like terms

OpenStudy (anonymous):

not quite the \(2x^2\) part is right, but \(2x\times (-x^2)=-2x^3\)

OpenStudy (anonymous):

multiplying like that has nothing to do with 'like terms' it is only when you add or subtract that you combine like terms

OpenStudy (anonymous):

so -x^3+(-2x^3)? or more?

OpenStudy (anonymous):

like terms?

OpenStudy (anonymous):

yes you can always multiply for example \(x\times x^2=x^3\) but when you add you can only add like terms for example \[x+3x=4x\] but you cannot add \(x+3x^2\)

OpenStudy (anonymous):

you know what i mean by like terms? same variable, same exponent

OpenStudy (anonymous):

lets do this from the beginning \[ -x^3+2x(x-x^2)\] multiply using the distributive property \[-x^3+2x\times x+2x\times (-x^2)\] \[-x^3+2x^2-2x^2\]

OpenStudy (anonymous):

then the "like terms" as \(-x^3\) and \(-2x^3\) but not the \(2x^2\)

OpenStudy (anonymous):

combining them gives \[-3x^3+2x^2\]

OpenStudy (anonymous):

Yeah, we learned like terms today

OpenStudy (anonymous):

it is clear if you think about adding numbers for example \[200=2\times 10^2,400=4\times 10^2\] and \[200+400=600=6\times 10^2\]

OpenStudy (anonymous):

so like for example, 5-x+2? How would that be done since it has no like terms? would I do stay change change and turn it to 5+(-x)+2 and then so 7+(-x)?

OpenStudy (anonymous):

like saying \(2x^2+4x^2=6x^2\)

OpenStudy (anonymous):

yes the numbers with no variables (constants) are also like terms, so of course you can add them or subtract them

OpenStudy (anonymous):

and your answer is correct, \(7+(-x)\) although you would probably write \(7-x\) same thing, just like \(10+(-4)=10-4\)

OpenStudy (anonymous):

ok I actually think I understand now

OpenStudy (anonymous):

you got another one we can try? or are you done

OpenStudy (anonymous):

I could check one, just one second

OpenStudy (anonymous):

k

OpenStudy (anonymous):

8b+5-3b

OpenStudy (anonymous):

So I would add 8b and 5? to make 13b-3b?

OpenStudy (anonymous):

no you cannot add \(8b\) and \(5\) since you don't know what \(b\) is but you can subtract \(8b-3b\)

OpenStudy (anonymous):

oh oops looked at that wrong

OpenStudy (anonymous):

suppose \(b\) as a \(\$100\) bill then you would have \(8\) \(\$100\) bills and then \(-3\) \($100\) bills leaving you with \(5\) \(\$100\) bills

OpenStudy (anonymous):

but that number \(5\) just stays there, because you have no other numbers to add or subtract

OpenStudy (anonymous):

so how would I simplify that? like step by step?

OpenStudy (anonymous):

\[8b+5-3b \] lets group the like terms together \[8b-5b+5\] now all that is left to do is find \(8b-3b\)

OpenStudy (anonymous):

hope it is clear that \(8b-3b=5b\) just like 8 million dollars minus 3 million dollars is 5 million dollars, and 8 bottle of beer minus 3 bottles of beer is 5 bottles of beer

OpenStudy (anonymous):

so final answer is \[5b+5\] that number 5 has no match, no other number to combine it with, so it just sits there

OpenStudy (anonymous):

ooookay now I get that one

OpenStudy (anonymous):

finally done with simplifying

OpenStudy (anonymous):

btw there is no such mathematical operation as "simplify" it is the lazy math teacher's way of saying "give me the answer i want to see" what you are doing is combining like terms

OpenStudy (anonymous):

hope you get it well, because it is kind of important

OpenStudy (anonymous):

laughing so hard at that bc its so true

OpenStudy (anonymous):

sometimes "simplify" means \[\frac{12}{8}=\frac{3}{2}\] which is really "reduce to lowest terms" you see that has nothing to do with what you are doing

OpenStudy (anonymous):

sometimes it means \(\sqrt{50}=5\sqrt{2}\) math teachers can say it to mean almost anything

OpenStudy (anonymous):

thank you sooooooo much

OpenStudy (anonymous):

yw